# Taking the Next Step with Generative Artificial Intelligence: The Transformative Role of Multimodal Large Language Models in Science Education

### v2

@@ 3. Applications of Multimodal LLMs for Science Education
The following chapter is structured by central aspects of science education (see 2.1). It aims to present exemplary potentials of MLLMs (see 2.3) in science teaching and learning, focusing specifically on adaptive, multimodal learning (see 2.4).
@@ 3.1. MLLMs for Content Creation
MLLMs can help teachers create tailored, multimodal learning materials to meet the diverse needs of the students. The development of learning materials is a central task for teachers. These materials should be designed to meet the learners’ needs while keeping the learning goal(s) in mind. Multiple representations can foster student motivation, interest, and conceptual understanding (Treagust,, 2008). For example, a diagram can provide learners with a way of visualizing the concept and hence developing a mental model (Gabel, 1998) (Gabel,, 1998). The creation of effective learning materials, therefore, often entails a shift in the modality - especially when the teacher has to cater materials to diverse learning groups.
While more skilled students might work on plain tabular data, others might be served better by providing a diagram or even (additional) textual or auditive explanations. Another example could be to provide accompanying visual information to otherwise text-only knowledge representations. This could be to map certain habitats, give images of listed laboratory equipment, provide images of certain species or flowers as well as provide structural animations of atoms and molecules. Using MLLMs, images can be flexibly adapted to make it easier for learners to directly recognize essential aspects of the visual representation (cf. signaling).
In addition, MLLMs can help organize learning content effectively, thereby reducing cognitive load, as (Mayer and Moreno,, 2003) suggested. The organization is crucial for creating materials that are accessible and understandable to a wide range of learners. MLLMs can promote active engagement with learning materials through generative activities, supporting (Mayer,, 2021) theory of integrating multimodal representations, shifting traditional content creation partially from the teacher to the student. In addition, MLLM-based learning content can be integrated into innovative, immersive virtual-reality learning environments, particularly due to their digital and spatial nature (see (Hartmann et al.,, 2023)).
@@ 3.2. MLLMs for Supporting and Empowering Learning
3.2.1. Fostering Scientific Content Knowledge
Engagement in real-world science materials is key to an authentic learning experience. Authentic learning can help boost motivation (Banas and York,, 2014) and science understanding (Nathan & Sawyer, 2014) . Learning with real-world materials, like research papers or even Wikipedia articles implies they were not originally designed (primarily) for learning, many original scientific sources are predominantly text-based.
MLLMs can enhance these materials by transforming or supplementing textual information with visual models, thereby aiding students in visualizing complex scientific concepts and increasing the accessibility of original sources. This gives students the opportunity to engage with real-world materials of their choice. For example, students try to understand the function of an insect’s eye based on an image they found on Wikipedia. MLLMs could help students to adaptively interact with these materials and complement text-based information (Figure Figure 1). The MLLM is capable of decoding the images information and providing adaptive information. Engagement in real-world science materials is key to an authentic learning experience. Authentic learning can help boost motivation (Banas and York,, 2014) and science understanding (Nathan & Sawyer, 2014). Learning with real-world materials, like research papers or even Wikipedia articles implies they were not originally designed (primarily) for learning, many original scientific sources are predominantly text-based.
Figure 1. Example of fostering scientific content knowledge: Graphical representation of an insects eye uploaded to an MLLM (ChatGPT mit GPT4-Vision, (OpenAI,, 2023)) asking for explanatory information for a 5th as well as a 12th grade student
Learners often have difficulties in understanding and interpreting diagrams (Henderson,, 1999). It is crucial to keep learners’ cognitive demands as low as necessary for efficient learning processes to take place (Mayer and Moreno,, 2003) and further to keep the text and image information understandable for the learner (cf. “pretraining principle”; e.g., (Jung et al.,, 2019)). MLLMs could potentially assist in the reception and understanding of diagrams, making complex data more accessible. This could be achieved by augmenting the diagram with textual explanations on the level of the diagram itself, like the explanation of the labels of the axis or even textualizing the presented data.
@@ 3.2.2. Fostering the Uses of Scientific Language
Learning the language of science is a major aspect of science education (Wellington and Osborne,, 2001; Nielsen,, 2012). But non-scientists, like students, often find scientific language to be severely limiting and difficult (Gardner,, 2012). From the perspective of multimedia learning, it is essential that the text and image information is understandable to the learner (Jung et al.,, 2019). MLLMs could be used to help students recall and remember scientific terms in images or diagrams. Using MLLMs, learners can adaptively simplify verbal information, i.e., technical language, or have technical terms explained.
@@ 3.2.3. Supporting Engagement in Scientific Practices
Engaging in scientific practices, such as modeling or planning and executing experiments, are central components of science curricula in Germany (KMK,, 2004) and around the world (e.g., US: National Research Council, 2012; UK: Department for Education, 2014; Finland: Finnish National Board of Education, 2014). MLLMs have the potential to assist students in various scientific practices, from generating ideas and models to interpreting diagrams and guiding investigations.
@@
When confronted with a scientific problem, students often struggle to formulate adequate research questions eligible for scientific investigation (Neber and Anton,, 2008; Hofstein et al.,, 2005). MLLMs can inspire students to think about potential scientific questions and formulate adequate hypotheses. For example, a teacher could provide an image of an eclipse, asking students to formulate questions they want to investigate. If some students struggle to develop adequate hypotheses on their own, the teacher has to provide appropriate scaffolds like providing more images, some complementary textual information, or motion graphics like videos and animations of the eclipse to prevent stagnation in the students’ scientific investigation. MLLMs could help generate appropriate scaffolds in various modalities on-the-fly. Equipped with tailored scaffolds, still remains up to the teacher to choose the appropriate scaffold. Visualizing and interpreting raw, unstructured data often challenges many students e.g., (von Kotzebue et al.,, 2014; Setiawan and Sukoco,, 2021).
@@
MLLMs possess the capability to render raw data into diagrams. Furthermore, MLLMs could significantly enhance understanding by explaining the key properties of the visualized data. These properties might include information about the x- and y-axes, underlying mathematical functions, or special points of interest within the data. MLLMs can provide tailored support by potentially shifting or adding another modality to the mathematical or data level, facilitating a deeper understanding. Moreover, they can assist in interpreting the data or drawing conclusions, thereby aiding in a comprehensive understanding of the data’s significance. By providing contextual explanations alongside visual representations, MLLMs can bridge the gap between raw data and meaningful interpretation, aiding not only in comprehension but also in deriving interpretations and conclusions from the data. Engaging with specialized laboratory equipment is an essential aspect of students’ scientific inquiry processes but instructions often refer to laboratory equipment in technical terms.
@@
Additionally, students sometimes also have difficulties with properly handling measuring and laboratory equipment (Kechel,, 2016). MLLMs can offer images of unknown instruments, preventing misuse and confusion. Beyond the static augmentation of names of laboratory equipment by explanatory images, MLLMs hold the potential to guide students during the use of laboratory equipment. By capturing photos of laboratory equipment with devices like smartphones or tablets and detailing the issues faced during operation, MLLMs can effectively guide students through the process, adapting responsively to their needs Figure 2.
@@
Note-taking is crucial for the success of scientific inquiry, yet it can be a complex task. To be successful, students first need a basic understanding of what and how to write down their notes (MacNeil and Falconer,, 2010), which draws attention during the process of scientific inquiry. MLLMs can help by transcribing spoken words into text summaries, which aids in note-taking without drawing any attention and cognitive capacities away from the scientific inquiry. The reduction of cognitive load might be beneficial, especially for introductory science courses. All these examples demonstrate the variety of how MLLMs can be employed in scientific practices, always aligning with basic multi-modal learning theories: multimodal aids can help students grasp complex scientific concepts and formulate relevant questions, which aligns with (Mayer,, 2021) and (Schnotz and Bannert,, 2003) theories on the integration of text and images, facilitating the construction of mental models from different representations. For instance, the assistance in the proper use of laboratory equipment by providing images and explanatory content aligns with the call for effective organization of learning materials by (Mayer and Moreno,, 2003). Transcription assistance is a practical application of the signaling principle, potentially reducing the cognitive load on students and aiding in the efficient organization of information.
@@ 3.2.4. Supporting Scientific Communication & Presentation
Communication of scientific results is a central aspect in science curricula (NRC, 2014; Department of Education, 2014; KMK, 2004). To be able to effectively communicate and present scientific ideas, students have to integrate new information with their existing knowledge to make sense of and create their content (Mayer,, 2021). This integration of multimodal representations requires generative activities on the part of the learner. MLLMs can convert data structures like tabular data into diagrams and animations or graphics into text, supporting the shift of modalities and, thus, the process of creating presentations. One promising example is provided in figure Figure 2 where a shift from tabular data to a diagram and, finally, textual data is demonstrated.
@@
Another example is the adaptive generation of complete image-based storyboards building on given analogies of scientific phenomena, as demonstrated by (Cao et al.,, 2023). In this process, firstly, abstract scientific concepts are translated into relatable text-based analogies to simplify understanding; secondly, these analogies are further converted into static visual representations to enhance conceptual clarity; and finally, these static visuals are evolved into dynamic visual analogies, offering dynamic representations.
@@ 3.3.1. Visual Assessment
Assessment is crucial for successful learning trajectories (Hattie,, 2008; Hattie and Timperley,, 2007). Especially forms of (computer-based) ‘rapid formative assessment’ have shown to be highly effective (Yeh,, 2010). There are already many tools that use LLMs to provide assessment in science education, e.g., (Bewersdorff et al.,, 2023; Zhai,, 2023). All of these mentioned assessment tools provide textual information based on textual input; they are uni-modal. Therefore, the ways of assessment are limited, especially in science education, with their predominant multimodal representations. MLLMs can deliver on any modality, enhancing the possibilities of assessment. For example, in problem-based learning scenarios, students’ self-generated and partially incorrect or incomplete problem-solving attempts can be used to prepare learners for future learning of mathematical concepts (cf. (Hartmann et al., 2022a, ); (Hartmann et al., 2022b, )). MLLMs can provide adaptive feedback and support to learners as they attempt to solve complex problems by automatically detecting learners’ problem-solving attempts. Furthermore, visual representations in science education encompass a wide variety of elements, including, but not limited to, mathematical graphs, diagrams, figures illustrating models and processes, as well as students’ sketches from experiments. They serve as various channels to convey and assimilate scientific ideas. However, there is a noticeable lack of using the full potential of assessing these visual representations. This primarily stems from the substantial time and effort required on the part of teachers to evaluate these often very complex, incomplete, or even contradictory visual constructs. Students often create written reports on an experiment, including drawings of their experimental setup. MLLMs, with their capability to assess text and visual content like graphs and diagrams, could provide a comprehensive, personalized assessment of their written reports, ultimately providing a deeper understanding of students’ cognitive processes and their skill in conveying complex science in both written and visual modes. Besides analyzing students’ cognitive processes by using MLLMs for assessment, teachers could potentially save time and enhance both the quality and objectivity of assessments.
@@ 3.3.2. Multimodal Feedback
Feedback, especially if elaborated and not only providing the correct answer, has been shown to be highly effective for students’ learning (Van der Kleij et al.,, 2015). Building on successful uni- or multimodal assessment, feedback can make student learning more effective. As with assessment, current feedback tools are mostly text-based e.g., (Seßler et al.,, 2023) and, therefore, might not use their full potential to provide effective feedback in science education. (Lee and Zhai,, 2023) demonstrated an AI system assessing drawn models of scientific phenomena. Their work paves the way for more comprehensive, multimodal feedback tailored to the student’s needs and the multimodal nature of science education (Unsworth, 2004) (Van Leeuwen and Kress,, 1990; Doran et al.,, 2021).
Students could receive feedback complemented by visual aids for their textual work. For example, visual representations like models of atoms, molecules, and their reactions could be provided to students describing a chemical reaction, helping them clarify misconceptions. Due to the responsiveness of MLLMs, these figures do not have to be stored and manually annotated to the particular misconceptions as in common, non-MLLM-based feedback tools. The design of feedback should always be considered in relation to subsequent related tasks (Henderson et al., 2019). MLLMs possess the capability to provide almost instant feedback, not only on student texts but also on other modalities like their drawings. This immediate response enables students to promptly adjust their work, thereby preventing them from pursuing ineffective approaches that could halt their learning process.

### v3

@@ 4. Applications of Multimodal LLMs for Science Education
The following section is structured by central aspects of science education (see 2.1). It aims to present exemplary potentials of MLLMs in science teaching and learning, focusing specifically on adaptive, multimodal learning. The given examples will be discusses in the light of our derived framework (see 3.). The examples discussed in the remainder of this section are succinctly represented in  Table 1.
Subsection Exemplary Aspects
3.1 MLLMs for Content Creation - Tailoring multimodal learning materials to diverse student needs - Organizing content effectively to reduce cognitive load - Promoting active engagement through generative activities - Flexible adaptation of visual representations for easier recognition (signaling) - MLLM-based code and content generation for virtual reality environments - Comprehensive integration of MLLMs into virtual reality settings via APIs
3.2 MLLMs for Supporting and Empowering Learning 3.2.1 Fostering Scientific Content Knowledge - Transforming/supplementing text with visuals 3.2.2 Fostering the Uses of Scientific Language - Assisting in understanding and using scientific language - Simplifying technical language and explaining technical terms 3.2.3 Supporting Engagement in Scientific Practices - Assisting in formulating research questions and hypotheses - Visualizing and interpreting raw data - Providing contextual explanations alongside visual representations. 3.2.4 Supporting Scientific Communication & Presentation - Converting data structures for effective communication - Generating image-based storyboards from analogies of scientific phenomena
3.3 MLLMs for Assessment and Feedback 3.3.1 Visual Assessment - Personalized assessment of text and visual content in students’ reports - Enhancing quality and objectivity of assessments 3.3.2 Multimodal Feedback - Providing elaborated feedback with visual aids - Instant feedback on various modalities like texts and drawings
Table 1. Exemplary aspects of the proposed framework for science education using MLLMs.
@@ 4.1. MLLMs for Content Creation
The development of learning materials is a central task for educators. These materials should be designed to meet the learners’ needs while keeping the learning goal(s) in mind. Multiple representations can foster student motivation, interest, and conceptual understanding (Treagust,, 2008). For example, a diagram can provide learners with a way of visualizing the concept and hence developing a mental model (Gabel,, 1998). The creation of effective learning materials, therefore, often entails a shift in the modality - especially when the educator has to cater materials to diverse learning groups. MLLMs can help educators create tailored, multimodal learning materials to meet the diverse needs of the students.
While more skilled students might work on plain tabular data, others might be served better with a diagram or even (additional) textual or auditive explanations. The educator could provide accompanying visual information to otherwise text-only knowledge representations. For example, this could be done to map certain habitats, give images of listed laboratory equipment, provide images of certain mammals or plants as well as provide structural visualisations of atoms and molecules. Using MLLMs, images can be flexibly adapted to make it easier for learners to directly recognize essential aspects of the visual representation (cf. signaling).
In addition, MLLMs can help organize learning content effectively, thereby reducing cognitive load, as Mayer and Moreno, (2003) suggested. The organization is crucial for creating materials that are accessible and understandable to a wide range of learners. For example teaching complex organic chemistry reactions, like the mechanism of SN1 and SN2 reactions, can be challenging due to their abstract nature. MLLMs could generate detailed, step-by-step visualizations of these mechanisms. Due to the adaptive nature of MLLMs, students can manipulate variables and see the effects on the reaction process.
Fostering engagement with such learning materials via MLLMs through generative activities, integrating multimodal representations (Mayer,, 2021), could even go further, shifting traditional content creation partially from the educator to the student: The student adaptively creates (organizes) his own learning materials based on an initial impulse, such as a textual description of the mechanism of SN1 and SN2 reactions - this can range from figures and diagrams to animations, or even synthetic datasets to investigate the chemical reaction.
Viewing the examples through the lens of the framework introduced in section 3, all the provided examples perform a shift from uni- to multimodality by adding a modality. One existing modality is extended with a second modality, potentially reducing cognitive load. In the examples given, the user of the MLLM is either the educator or the student themselves. Both can create adaptive learning materials.
Moreover, MLLM-based learning content can be integrated into innovative, immersive virtual-reality learning environments. The ability of MLLMs to reason about spatial relationships, coupled with their ability to generate relevant code and visual elements, facilitates their use in the design, creation, and extension of virtual environments. In addition, well-defined application programming interfaces (APIs) to MLLMs allow for a comprehensive integration of MLLMs into these virtual settings (see Hartmann et al., (2023)).
@@ 4.2. MLLMs for Supporting and Empowering Learning
4.2.1. Fostering Scientific Content Knowledge Construction
Engagement in real-world science materials is key to an authentic learning experience. Authentic learning can help boost motivation (Banas and York,, 2014) and science understanding (Nathan and Sawyer,, 2014). Learning with real-world materials, like research papers or even Wikipedia articles, implies they were not originally designed (primarily) for learning; many original scientific sources are predominantly text-based.
MLLMs can enhance these materials by supplementing textual information with visual models (see Chapter 3), thereby aiding students in visualizing complex scientific concepts and increasing the accessibility of original sources. This gives students the opportunity to engage with real-world materials of their choice. For example, students try to understand the function of an insect’s compound eye based on an image they found on Wikipedia. MLLMs could help students to adaptively interact with these materials and complement text-based information (Figure 2) based on their background and prior knowledge. As exemplarily depicted in Figure 2, the MLLM is capable of analyzing the image and providing adaptive information tailored to the competency of the learner (i.e., grade 5 and grade 12 student).
Figure 2. Example of fostering scientific content knowledge: Graphical representation of an insect’s eye uploaded to an MLLM (ChatGPT with GPT4-Vision, (OpenAI,, 2023)) asking for explanatory information for a 5th as well as a 12th grade student.
It is well known in previous works in education research that learners often face difficulties in understanding and interpreting diagrams (Henderson,, 1999). Therefore, it is crucial to keep learners’ cognitive demands as low as necessary for efficient learning processes to take place (Mayer and Moreno,, 2003) and further to keep the text and image information understandable for the learner (cf. “pretraining principle”; e.g., Jung et al., (2019)). In this context, MLLMs could assist in the reception and understanding of diagrams, making complex data more accessible. This could be achieved by augmenting the diagram with textual explanations on the level of the diagram itself, like the explanation of the labels of the axis or even textualizing the presented data.
Both examples, when viewed through the framework introduced in Section 3, demonstrate a shift from unimodality to multimodality by incorporating an additional modality. The visual modality, which includes a graphical representation of an insect’s eye or a diagram, is augmented with textual information. The user of the MLLM can be in both examples either the learner or the educator.
@@ 4.2.2. Fostering the Uses of Scientific Language
Learning scientific language is a major aspect of science education (Wellington and Osborne,, 2001; Nielsen,, 2012). But non-scientists, like students, often find scientific language to be severely limiting and difficult (Gardner,, 2012). From the perspective of multimedia learning, it is essential that the text and image information is understandable to the learner (Jung et al.,, 2019). MLLMs could be used in this context to help students recall and remember scientific terms in images or diagrams. Using MLLMs, learners can adaptively augment textual information like technical language, or have technical terms explained on the fly, for example, from a scientific podcast. Previous unimodal learning material, visuals such as images or diagrams, are supplemented by a second modality, textual information like technical language (cf. Section 3).
@@ 4.2.3. Supporting Engagement in Scientific Practices
Engaging in scientific practices, such as modeling or planning and executing experiments, are central components of science curricula in Germany (KMK,, 2004) and around the world (e.g., US: NRC, (2012), UK: Department for Education, (2013), and globally: OECD, (2018)). MLLMs have the potential to assist students in various scientific practices, from generating ideas and models to interpreting diagrams and guiding investigations.
@@
When confronted with a scientific problem, students often struggle to formulate adequate research questions eligible for scientific investigation (Neber and Anton,, 2008; Hofstein et al.,, 2005). MLLMs can inspire students to think about potential scientific questions and help in the process of formulating adequate hypotheses. For example, an educator could provide an image of an eclipse, asking students to formulate questions they want to investigate. If some students struggle to develop adequate hypotheses on their own, the educator has to provide appropriate scaffolds like providing more images, some complementary textual information, or motion graphics like videos and animations of the eclipse to prevent stagnation in the students’ scientific investigation. MLLMs could help generate appropriate scaffolds in different modalities on-the-fly. Equipped with tailored scaffolds, it still remains up to the educator to choose the appropriate degree of guidance.
@@
Furthermore, visualizing and interpreting raw, unstructured data often challenges many students (e.g., von Kotzebue et al., (2014); Setiawan and Sukoco, (2021)). MLLMs have the capability to render raw data into diagrams. Furthermore, MLLMs could significantly enhance understanding by explaining the key properties of the visualized data. These properties might include information about the x- and y-axes, underlying mathematical functions, or special points of interest within the data. MLLMs can provide tailored support by potentially shifting or adding another modality to the mathematical or data level, facilitating thus a deeper understanding. Moreover, they can assist in interpreting the data or drawing conclusions, thereby aiding in a comprehensive understanding of the data’s significance. By providing contextual explanations alongside visual representations, MLLMs can bridge the gap between raw data and meaningful interpretation, aiding not only in comprehension but also in deriving interpretations and conclusions from the data.
@@
Engaging with specialized laboratory equipment is an essential aspect of students’ scientific inquiry processes, but instructions often refer to laboratory equipment in technical terms. Additionally, students sometimes also have difficulties with properly handling measuring and laboratory equipment (Kechel,, 2016). In this context, MLLMs can offer images of unknown instruments, preventing thus misuse and confusion. Beyond the static augmentation of names of laboratory equipment by explanatory images, MLLMs hold the potential to guide students during the use of laboratory equipment. By capturing photos of laboratory equipment with devices like smartphones or tablets and detailing the issues faced during operation, MLLMs can effectively guide students through the process, adapting responsively to their needs (see Figure 3).
@@
Note-taking is crucial for the success of scientific inquiry, yet it can be a complex task. To be successful, students first need a basic understanding of what and how to write down their notes (MacNeil and Falconer,, 2010), which draws attention and ultimatively cognitive load during the process of scientific inquiry. MLLMs can help by transcribing spoken words into text summaries, which aids in note-taking without drawing any attention and cognitive capacities away from the scientific inquiry. Educators must always carefully consider how to provide scaffolding in their teaching without leading students to, at worst, mindlessly follow procedures and outsource the actual thinking to the MLLM. Reducing cognitive load can be especially beneficial for introductory science courses.
All these examples demonstrate the variety of how MLLMs can be employed in scientific practices, always aligning with basic multi-modal learning theories: multimodal aids can help students grasp complex scientific concepts and formulate relevant questions, which aligns with Mayer, (2021) and Schnotz and Bannert, (2003) theories on the integration of text and images, facilitating the construction of mental models from different representations. In our framework (Section 3), this process relates to augmenting one modality with a second modality. The user can be either the educator or the student. Based on information provided by the user, the MLLM offers adaptive learning materials in a second modality.
@@ 4.2.4. Supporting Scientific Communication & Presentation
Communication of scientific results and ideas is a central aspect in science curricula (Department for Education,, 2013; NRC,, 2012; KMK,, 2004). For students to proficiently convey and articulate scientific concepts, they must generate new content. Therefore, it is important that they actively integrate new information with their pre-existing knowledge to make sense of and create their content (Mayer,, 2021). This integration of multimodal representations requires generative activities like creating diagrams from data by the student. MLLMs can assist in this process by converting data structures, such as tabular data, into diagrams, animations, or graphics, and transforming graphics into text, thereby supporting the process of creating content for scientific communication. One example is provided in Figure 4 where a shift from tabular data to a diagram and, finally, textual data is demonstrated.
@@
Another example is the adaptive generation of complete image-based storyboards building on given analogies of scientific phenomena, as demonstrated by Cao et al., (2023) (example: ibidem). In this process, firstly, abstract scientific concepts are translated into relatable text-based analogies to simplify understanding; secondly, these analogies are further converted into static visual representations to enhance conceptual clarity; and finally, these static visuals are evolved into dynamic visual analogies, offering dynamic representations.
In our framework (Section 3), the creation of diagrams and then generating explanatory text based on tabular data relates to transforming content from one modality to a second modality as well as augmenting one modality with another. Tabular data (textual) is transformed into a visual representation (diagram). This shift in modalities provides further insights into the data. In the second step, the diagram is augmented by textual information describing the diagram. Based on information provided by the user, the MLLM offers adaptive learning materials in a second modality.
@@ 4.3.1. Visual Assessment
Assessment is crucial for successful learning trajectories (Hattie,, 2008; Hattie and Timperley,, 2007). Especially forms of (computer-based) ‘rapid formative assessment’ have shown to be highly effective (Yeh,, 2010).
Visual representations in science education encompass a wide variety of elements, including, but not limited to, mathematical graphs, diagrams, figures illustrating models and processes, as well as students’ sketches from experiments. They serve as various channels to convey and assimilate scientific ideas. However, there is a noticeable lack of using the full potential of assessing these visual representations. This primarily stems from the substantial time and effort required on the part of educators to evaluate these often very complex, incomplete, or even contradictory visual constructs.
There are already many tools that use LLMs to provide assessment in science education, e.g., Bewersdorff et al., 2023a ; Zhai, (2023); Latif and Zhai, (2024). All of these mentioned assessment tools provide textual information based on textual input; they are uni-modal. Therefore, the ways of assessment are limited, especially in science education, with their predominant multimodal representations. MLLMs can deliver on any modality, enhancing the possibilities of assessment.
Students often create written reports on an experiment, including drawings of their experimental setup. MLLMs, with their capability to assess text and visual content like graphs and diagrams, could provide a comprehensive, personalized assessment of their written reports, ultimately providing a deeper understanding of students’ cognitive processes and their skill in conveying complex science in both written and visual modes. Besides analyzing students’ cognitive processes by using MLLMs for assessment, educators could potentially save time and enhance both the quality and objectivity of assessments. Lee and Zhai, (2023) demonstrated an AI system assessing drawn models of scientific phenomena, pioneering new ways of visual assessment. For visual assessment, the visual input, such as a student’s drawing, is transformed into a textual representation, such as written feedback or a written assessment (cf. section 3).
@@ 4.3.2. Multimodal Feedback
Feedback, especially if elaborated and not only providing the correct answer, has been shown to be highly effective for students’ learning (Van der Kleij et al.,, 2015). Building on successful uni- or multimodal assessment, feedback can make student learning more effective. As with assessment, current feedback tools are mostly text-based e.g., Seßler et al., (2023) and, therefore, might not use their full potential to provide effective feedback in science education. With regard to CTML (Mayer,, 2021), multimodal feedback, combining visual and textual elements, engages ideally dual-channel cognitive processing of learners. This integration not only balances cognitive load but also enhances comprehension and retention, ultimately fostering understanding.
Students could receive feedback complemented by visual aids for their textual work. For example, visual representations like models of atoms, molecules, and their reactions could be provided to students describing a chemical reaction, helping them clarify misconceptions. Due to the responsiveness of MLLMs, these figures do not have to be stored and manually annotated to the particular misconceptions as in common, non-MLLM-based feedback tools.
The design of feedback should always be considered in relation to subsequent related tasks (Henderson et al.,, 2019). MLLMs possess the capability to provide almost instant feedback, not only on student texts but also on other modalities like their drawings. This immediate response enables students to promptly adjust their work, thereby preventing them from pursuing ineffective approaches that could halt their learning process. For multimodal feedback, traditional textual feedback can be augmented with an additional visual modality like a diagram. The user of the MLLM can be either the educator or the student, as feedback may be directed to the student for personal reflection or to the educator for further refinement or validation. Regardless of the recipient, feedback can always be personalized and adaptive by the MLLM (cf. section 3).